OpenStudy (anonymous):

Which expression is equivalent to: sin^2 x -sin x-2/sin x-2? a)sin x+1 b)sin x-1 c)sin^2x d)-sin^2x Could anyone explain to me how to do this?

OpenStudy (anonymous):

sin^2 x -sin x-2/(sin x-2) sin^2 x -2sin x +sinx -2/(sin x-2) sinx(sinx -2) +1(sinx -2)/(sinx-2) =(sinx-2)(sinx+1)/(sinx-2) =sinx +1

OpenStudy (alekos):

is the last term (sin(x-2))/(sin(x-2)) ?

OpenStudy (anonymous):

Thank you so much!! @matricked

OpenStudy (anonymous):

yw